A passenger train and a freight train start at the same point at the same time and travel in opposite directions. The passenger train travels twice as fast as the freight train. In 4 hours they are 396 miles apart. What is the rate of each train
Answer by askmemath(368)

Let the speed of the 1st train = X
Then speed of second = 2X
Each travels for 4 hours
So distance covered by each = 4X+8X = 12X
But this is given as 396
So 12X = 396
X = 396/12 = 35.5mph
Speed of the second train = 2 x 35.5 = 71mph